OpenStudy (anonymous):

Medal Find the product of 5/6 , 9/32 and 24/35 Express your answer as a fully reduced fraction.

OpenStudy (danjs):

\[\frac{ 5*9*24 }{ 6*32*35 }\]

OpenStudy (anonymous):

164.57

ganeshie8 (ganeshie8):

you want answer as a fraction (like a/b) rihgt ? how did you get 164.57 ?

OpenStudy (anonymous):

5*9*24 / 6*32*35

OpenStudy (danjs):

yes, now you have to reduce that to lowest terms

OpenStudy (danjs):

i would divide both top and bottom by 6 to begin with

OpenStudy (anonymous):

1080/ 6720

OpenStudy (danjs):

it is easier to leave it as you had it, more simple to see the common factors to reduce the fraction

OpenStudy (danjs):

as you can see the top has 24 and the bottom has a 6, so you can divide both top and bottom by 6

OpenStudy (anonymous):

3/336

OpenStudy (danjs):

\[\frac{ 5*9*24 }{ 6*32*35} = \frac{ 5 * 9 * 4 }{ 1 * 32 * 35 }\] see how both top and bottom were divided by 6 to reduce the fraction?

OpenStudy (anonymous):

180/1120

OpenStudy (danjs):

so, now you see you have a 5 in the top and a 35 in the bottom, so you can divide both those by 5

OpenStudy (anonymous):

36/224

OpenStudy (danjs):

\[\frac{ 1*9*4 }{ 1*32*7 }\]

OpenStudy (danjs):

do not calculate, to reduce a fraction you need to know the common factors of the numerator and the denominator

OpenStudy (danjs):

What is the next step from the last one i wrote?

OpenStudy (anonymous):

36/224 = 18/112 = 9/56

OpenStudy (anonymous):

your hot@harz360

OpenStudy (danjs):

\[\frac{ 1*9*4 }{ 1*32*7 } = \frac{ 1*9*1 }{ 1*8*7 }\] I divided top and bottom by 4

OpenStudy (anonymous):

9/56

OpenStudy (danjs):

yes, but you need to know how to do it, for instance if you have ... 6588 / 10589 how are you going to reduce that into lowest terms? you need to factor out common factors as in the process i went through.

OpenStudy (anonymous):

so the answer is 9/56?

OpenStudy (danjs):

unless you can divide both 9 and 56 by a common factor to reduce it further.

OpenStudy (anonymous):

u cant reduce 9/56

OpenStudy (danjs):

correct

OpenStudy (anonymous):

so the answer 9/56?